OpenStudy (anonymous):

A nick on the edge of a CD rotates to (−6, 5) during one song when represented graphically. What is the sine value of this function?

OpenStudy (nurali):

The -6 and the 5 are the legs of a right triangle. The hypotenuse can be found bt the pythagorean theorem: H = sqrt((-6)^2 + 5^2) = 7.810 The value of the sine at that point would be: sin(theta) = 5/7.81 = 0.6042

OpenStudy (nurali):

Welcome to Openstudy.